OpenStudy (anonymous):

find the products of the polynomials below. enter your answer in descending order in the box below enter exponents using the caret (^). (x^3+2x+2)(x^2-x-1)

OpenStudy (anonymous):

\[x^{a} \times x^{b} = x^{a+b}\] first one is x^3 times x^2 = x^5

OpenStudy (zale101):

@nincompoop